OpenStudy (anonymous):

Determine the values of x for which f'(x)=0, given that f(x)=2x/(1+4x^2).

OpenStudy (anonymous):

Derivative is (-8x^2+2)/(1+4x^2)^2.

OpenStudy (anonymous):

then set the numerator equal to zero and solve

OpenStudy (anonymous):

I know, but I'm getting confused on how to separate x and whatnot.

OpenStudy (anonymous):

\[8x^2=2\] \[x^2=\frac{1}{4}\] \[x=\pm\frac{1}{2}\]

OpenStudy (anonymous):

Oh, numerator only you said.

OpenStudy (anonymous):

yea, fraction is only zero when numerator is

OpenStudy (anonymous):

Is that always the case when you are looking for a value of a fraction that equals zero?

OpenStudy (anonymous):

sure, just like with numbers right?

OpenStudy (anonymous):

an x value that is

OpenStudy (anonymous):

\[\frac{0}{b}=0\]

OpenStudy (anonymous):

Yep, true. Ok makes sense. Thanks.